A car accelerates uniformly from rest at 4m/s². How far will it travel in the fifth second?
1
Expert's answer
2020-07-16T10:58:15-0400

"S_5=\frac{at_5^2}{2}=\frac{4\cdot5^2}{2}=50 m"


"S_4=\frac{at_4^2}{2}=\frac{4\cdot4^2}{2}=32m"


Distance covered in "5^{th}" second


"S_5-S_4=50-32=18m" Answer